### v2

# RGBSQGrasp: Inferring Local Superquadric Primitives from Single RGB image for Graspability-Aware Bin Picking

Abstract
Superquadrics (SQ) offer a compact, interpretable shape representation that captures the physical and graspability understanding of objects. In this work, we propose RGBSQGrasp, a grasping framework that leverages superquadric shape primitives and foundation metric depth estimation models to infer grasp poses from a monocular RGB camera—eliminating the need for depth sensors. Our framework integrates a cross-platform dataset generation pipeline, a foundation model-based object point cloud estimation module, a global-local superquadric fitting network, and an SQ-guided grasp pose sampling module. By integrating these components, RGBSQGrasp reliably infers grasp poses through geometric reasoning. Real-world robotic experiments demonstrate a 92% grasp success rate, highlighting the effectiveness of RGBSQGrasp in packed bin-picking environments. Supplementary demonstration videos can be found at project website: https://rgbsqgrasp.github.io/
Figure 2: We illustrate a sequential rollout of the superquadrics-guided robotic grasping process. The red point cloud represents the superquadric fitting for each partial point cloud, while the green vector denotes the grasp sampled from the fitted superquadrics.
I INTRODUCTION
In industrial environments such as warehousing and manufacturing, robotic bin picking is crucial for tasks like inventory management and assembly line operations [1]. Bin environments present two major challenges: (1) the available camera views are limited, and (2) tight spatial constraints and occlusions caused by object-to-object and object-to-bin interactions [2]. To address these issues, superquadric primitives have gained increasing interest in robotic grasping due to their compact parametrization and generalization to novel shapes, offering advantages in both object understanding and grasp evaluation [3, 4, 5, 6, 7]. Superquadric fitting is typically framed as an optimization problem [8], requiring nearly complete point clouds input, which is impractical in settings with limited camera views. Previous work has explored using depth images for primitive recognition due to their richer geometric information [9, 10]. However, these methods often depend on high-quality depth sensors, as noise and discrepancies from real-world sensors degrade performance [10], thereby introducing expensive resource dependencies [11].
To address these challenges, we propose a superquadric-guided grasping framework for bin picking using a single RGB sensor. The framework comprises a universal, cross-platform synthetic dataset generation pipeline using physics engines, a data-driven superquadric fitting network to infer superquadrics from locally visible geometry, and an inference module that leverages state-of-the-art vision foundation models for enhanced generalization and robustness [12, 13]. Our key contributions are as follows:
  • We establish a superquadric-guided framework using a single monocular RGB image to infer superquadric representations without requiring prior object knowledge or multi-camera views.
  • This approach achieves a 92% success rate in packed bin-picking scenes by leveraging superquadric primitives for efficient grasp synthesis and evaluation.
II RELATED WORK
II-A Preliminary
Superquadrics are a versatile family of parametric geometric primitives that can represent a wide range of shapes, within a continuous parameter space [14]. They are commonly defined using an implicit function, as shown in Equation  1:
$F(\mathbf{x})=\left(\left(\frac{x}{a_{x}}\right)^{\frac{2}{\epsilon_{2}}}+\left(\frac{y}{a_{y}}\right)^{\frac{2}{\epsilon_{2}}}\right)^{\frac{\epsilon_{2}}{\epsilon_{1}}}+\left(\frac{z}{a_{z}}\right)^{\frac{2}{\epsilon_{1}}}-1,$ (1)
where $\mathbf{x}=[x,y,z]^{T}\in\mathbb{R}^{3}$ is a point in the superquadric’s local coordinate frame. The parameters $a_{x},a_{y},a_{z}>0$ represent scale factors along the $x$, $y$, and $z$-axes, respectively. The shape parameters $\epsilon_{1}$ and $\epsilon_{2}$ ($\epsilon_{1},\epsilon_{2}\geq 0$) control the curvature and convexity of the primitive.
III Methodology
Our method addresses the challenge of enabling robots to understand cluttered scenes and identify graspable regions from single monocular RGB input in bin-picking tasks, as illustrated by demonstration in Figure 2.
III-A Dataset Generation
The goal of our dataset generation is to create a diverse synthetic dataset. We constructed a synthetic primitive dataset by capturing partially observed point clouds for objects with varying occlusion levels in each scene. For each object and its partial point cloud (PCD), the ground truth superquadric (SQ) primitive is sampled from the object model, as shown in Figure LABEL:fig:RGBSQGrasp. Our synthetic dataset includes 36K PCD-SQ ground truth pairs for training.
III-B Superquadrics Fitting Network
Superquadrics Fitting model is trained to predict superquadric shape parameters from partial point clouds. The ground truth is generated by uniformly sampling the superquadric surface, generated from ground truth SQ parameters. The network architecture integrates both local and global feature extraction. The local feature extractor utilizes the DGCNN architecture [15]. The global feature extractor, adapted from PointNet [16], predicts scale parameters to capture the object’s overall geometry. For translation and rotation, the model computes a weighted ensemble of the outputs from both the global and local branches. The model uses four prediction heads: the shape head outputs $\epsilon_{1}$ and $\epsilon_{2}$, the scale head outputs $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$, the rotation head outputs a rotation matrix $\mathbf{R}$, and the translation head outputs a translation vector $\mathbf{t}$.
III-C SQ-guided Scene Understanding
For real-robot bin picking deployment, we utilize a Microsoft Azure Kinect RGB camera as a monocular sensor. We leverage the Depth Pro [17] foundation model for metric depth estimation and Segment Anything 2 [18] for depth-based segmentation. By projecting the metric depth onto the segmentation map, we obtain instance-aware PCD map. The preprocessed PCD is then passed through the trained superquadric fitting network to obtain the SQ parameters. Finally, the SQ point cloud is projected back into the scene for grasp sampling, with each SQ primitive capturing sufficient geometric information for the downstream grasping task.
III-D SQ-guided Grasp Pose Sampling
The sampling strategy follows the approach proposed in [19]. Key steps include sampling antipodal grasp poses and moving the gripper along specific axes, such as along the principal axes (x, y, and z), to generate a diverse set of candidate poses for each superquadric by leveraging the superquadric’s symmetry and geometric properties. After generating a set of potential grasp candidates, the grasp pose is further selected based on the SQ Quality Score. After aligning a partial object point cloud $S$ (source) with the corresponding SQ fitting $T$ (target), the SQ quality score $C_{i}$ for each target point $t_{i}$ is calculated based on its distance to the nearest neighbor source point $s_{j}$. If $\bar{C}_{k}>\sigma$, the grasp pose from that region is considered high-quality fitting region with valid SQ fitting.
IV Results and Future works
In this work, we propose a superquadric-guided framework for bin picking using a single monocular RGB image, enabling robust grasp synthesis without object priors. Real-robot experiments demonstrate its effectiveness, achieving a 92% grasp success rate over 189 picks; Some demonstration samples can be found here: https://www.youtube.com/watch?v=79vJCNBJ6-0. Future work will explore extending the framework to other end-effectors, such as suction cups and robotic hands, as the graspability analysis enabled by superquadrics is generalizable.